When we perform subtraction on -7 and 1 the answer in 2 s compliment form is _________

a) 1010

b) 1110

c) 0110

d) 1000

In radix binary, 1101 integer multipled by 1011 is

a) 10001111

b) 10101010

c) 11110000

d) 11001100
Answer #1
1)
-7 - 1 = -8
-8 in 2's complement is 1000
Answer: d) 1000


2)
1101 is 13
1011 is 11
13 * 11 = 143
143 in binary is 10001111
Answer: a) 10001111